# VGGT-SLAM: Dense RGB SLAM Optimized on the SL(4) Manifold

### v1

@@ 2 Related Work
To extend the idea of DUSt3R to multiple frames, Spann3R [67] leverages a learned memory module and Cut3R [69] uses a recurrent state model. Both can incrementally reconstruct a scene using multiple images, but are each limited to short sequences. Recently, Pow3R [27] extends the DUSt3R framework to optionally take in any estimates of any combination of camera intrinsics, poses, and depth (which may be sparse or dense) and demonstrates substantial improvement in scene reconstruction and pose estimation given the added inputs. Splatt3R [60] extends the DUSt3R idea to Gaussian Splatting [29] by directly outputting the the Gaussian Splatting parameters given two views and PreF3R [8] extends this to multiple views using a similar memory framework as Spann3R. Reloc3r [13] modifies the DUSt3R framework for directly outputs relative camera poses and uses motion averaging to recover absolute poses with respect to a map database.
@@ 4.4 Backend: Nonlinear factor graph optimization on the $\mathrm{SL}(4)$ manifold
To solve (3), we iteratively compute state increments by solving a linearized least squares problem. To this end, we define ${{\bm{\xi}}}\in\mathbb{R}^{15}$ as the tangent-space parameterization of $\mathrm{SL}(4)$, the mapping function $\mathrm{Exp}:\mathbb{R}^{15}\rightarrow\mathrm{SL}(4)$, which satisfies $\mathrm{Log}\big{(}\mathrm{Exp}\left({{\bm{\xi}}}\right)\big{)}={{\bm{\xi}}}$ and $\mathrm{Exp}\left({{\bm{\xi}}}\right)=\mathrm{exp}({{\bm{\xi}}}^{\wedge})=\mathbf{H}$. In particular, ${{\bm{\xi}}}^{\wedge}$ is a Lie algebra element of $\mathfrak{sl}(4)$, computed by summing the $k$-th component of ${{\bm{\xi}}}$ with its $k$-th corresponding generator $\mathbf{G}_{k}\;\forall k:\{1:15\}$ (i.e., ${{\bm{\xi}}}^{\wedge}=\sum_{k=1}^{15}{{\bm{\xi}}}_{k}\,\mathbf{G}_{k}$[15]. More details can be found in the supplementary materials.

### v2

@@ 2 Related Work
To extend the idea of DUSt3R to multiple frames, Spann3R [67] leverages a learned memory module and Cut3R [69] uses a recurrent state model. Both can incrementally reconstruct a scene using multiple images, but are each limited to short sequences. Recently, Pow3R [27] extends the DUSt3R framework to optionally take in any estimates of any combination of camera intrinsics, poses, and depth (which may be sparse or dense) and demonstrates substantial improvement in scene reconstruction and pose estimation given the added inputs. Splatt3R [60] extends the DUSt3R idea to Gaussian Splatting [29] by directly outputting the Gaussian Splatting parameters given two views, and PreF3R [8] extends this to multiple views using a similar memory framework as Spann3R. Reloc3r [13] modifies the DUSt3R framework for directly outputs relative camera poses and uses motion averaging to recover absolute poses with respect to a map database.
@@ 4.4 Backend: Nonlinear factor graph optimization on the $\mathrm{SL}(4)$ manifold
To solve (3), we iteratively compute state increments by solving a linearized least squares problem. To this end, we define ${{\bm{\xi}}}\in\mathbb{R}^{15}$ as the tangent-space parameterization of $\mathrm{SL}(4)$, the mapping function $\mathrm{Exp}:\mathbb{R}^{15}\rightarrow\mathrm{SL}(4)$, which satisfies $\mathrm{Log}\big{(}\mathrm{Exp}\left({{\bm{\xi}}}\right)\big{)}={{\bm{\xi}}}$ and $\mathrm{Exp}\left({{\bm{\xi}}}\right)=\mathrm{exp}({{\bm{\xi}}}^{\wedge})=\mathbf{H}$. In particular, ${{\bm{\xi}}}^{\wedge}$ is a Lie algebra element of $\mathfrak{sl}(4)$, computed by summing the $k$-th component of ${{\bm{\xi}}}$ with its $k$-th corresponding generator $\mathbf{G}_{k}\;\forall k:\{1:15\}$ (i.e., ${{\bm{\xi}}}^{\wedge}=\sum_{k=1}^{15}{{\bm{\xi}}}_{k}\,\mathbf{G}_{k}$[15]. More details can be found in Appendix A.
@@ 7 Conclusion
Acknowledgments
This work is supported in part by the NSF Graduate Research Fellowship Program under Grant 2141064, the ONR RAPID program, and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00461409). The authors would like to gratefully acknowledge Riku Murai for assisting us with benchmarking.
Appendix A Tangent space of $\mathrm{SL}(4)$
Here, we provide the explicit 15 generators, $\mathbf{G}_{k}\;\forall k:\{1:15\}$, of $\mathrm{SL}(4)$, which allow us to relate the Lie algebra $\mathfrak{sl}(4)$ to the Lie group $\mathrm{SL}(4)$.
The tangent space of $\mathrm{SL}(4)$ consists of all $4\times 4$ real matrices with zero trace. Thus, there are 15 generators, $\mathbf{G}_{k}$, where 12 of them are defined as ${\bm{E}}_{ab}$ for $a\neq b$ where 1 is in the $(a,b)$ entry and 0, elsewhere. The remaining three generators are ${\bm{B}}_{1}=\mathrm{diag}(1,-1,0,0)$, ${\bm{B}}_{2}=\mathrm{diag}(0,1,-1,0)$, ${\bm{B}}_{3}=\mathrm{diag}(0,0,1,-1)$. Explicitly, the generators are as follows:
$\mathbf{G}_{1}={\bm{E}}_{01}$ $=\begin{pmatrix}0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{2}={\bm{E}}_{02}$ $=\begin{pmatrix}0&0&1&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{3}={\bm{E}}_{03}$ $=\begin{pmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$
$\mathbf{G}_{4}={\bm{E}}_{10}$ $=\begin{pmatrix}0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{5}={\bm{E}}_{12}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{6}={\bm{E}}_{13}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$
$\mathbf{G}_{7}={\bm{E}}_{20}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{8}={\bm{E}}_{21}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{9}={\bm{E}}_{23}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\end{pmatrix},$
$\mathbf{G}_{10}={\bm{E}}_{30}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\end{pmatrix},$ $\mathbf{G}_{11}={\bm{E}}_{31}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&1&0&0\end{pmatrix},$ $\mathbf{G}_{12}={\bm{E}}_{32}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\end{pmatrix},$
$\mathbf{G}_{13}={\bm{B}}_{1}$ $=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{14}={\bm{B}}_{2}$ $=\begin{pmatrix}0&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&0\end{pmatrix},$ $\mathbf{G}_{15}={\bm{B}}_{3}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}.$
Thus, as briefly explained in Sec. 4.4, the relation between the Lie algebra, ${{\bm{\xi}}}^{\wedge}\in\mathfrak{sl}(4)$, and the Lie group $\mathbf{H}\in\mathrm{SL}(4)$ is given by:
$\mathbf{H}=\exp\left({{\bm{\xi}}}^{\wedge}\right)=\exp\left(\sum_{k=1}^{15}{{\bm{\xi}}}_{k}\,\mathbf{G}_{k}\right).$ (6)
Appendix B Extra Quantitative Results
We provide addition results of evaluating on the 7-Scenes [58] and TUM RGB-D [63] datasets where we experiment with different submap sizes (Sec. B.1) and show the number of submaps and loop closures per scene (Sec. B.2).
B.1 Evaluation with different submap sizes
Here we show results for the 7-Scenes and TUM RGB-D datasets in Tables 4 and 5 with different submap sizes ($w=8,16,32$). For 7-Scenes, we also include results for $w=1$. Recall that $w$ is the size of new frames in the submap, so in the case of $w=1$, each submap has one new frame, one frame from the prior submap, and up to one extra frame from loop closures. For small submap size of $w=1$, the backend becomes numerically unstable for some TUM scenes (consistently floor and 360) preventing an estimated alignment, and thus we do not include the $w=1$ for TUM. This is due to reasons discussed in Sec. 6. Particularly, for the floor scene there are a large portion of images which only view a planar scene which makes the estimation of the full 15-DOF homography matrix degenerate, and for the 360 scene, using a small submap size such as $w=1$ is likely to encounter a pure rotation which can result in less accurate depth measurements from VGGT and hence reduced accuracy in our estimate of relative homography.
Table 4: Root mean square error (RMSE) of absolute trajectory error (ATE) on 7-Scenes [58] (unit: m). The * symbol indicates that the baseline is evaluated in the uncalibrated mode, all VGGT-SLAM configurations are evaluated in uncalibrated mode. Green is best and light green is second best.
Method Sequence Avg
chess fire heads office pumpkin kitchen stairs
MASt3R-SLAM* [44] 0.063 0.046 0.029 0.103 0.114 0.074 0.032 0.066
Ours ($\mathrm{Sim}(3),w=1$) 0.047 0.025 0.032 0.113 0.138 0.050 0.083 0.070
Ours ($\mathrm{Sim}(3),w=8$) 0.039 0.027 0.020 0.108 0.144 0.053 0.080 0.067
Ours ($\mathrm{Sim}(3),w=16$) 0.037 0.027 0.021 0.107 0.135 0.051 0.093 0.067
Ours ($\mathrm{SL}(4),w=1$) 0.089 0.046 0.072 0.119 0.147 0.055 0.100 0.090
Ours ($\mathrm{SL}(4),w=8$) 0.041 0.060 0.043 0.106 0.206 0.054 0.078 0.084
Ours ($\mathrm{SL}(4),w=16$) 0.036 0.065 0.037 0.107 0.139 0.050 0.093 0.075
Table 5: Root mean square error (RMSE) of absolute trajectory error (ATE) on TUM RGB-D [63] (unit: m). The * symbol indicates that the baseline is evaluated in the uncalibrated mode, all VGGT-SLAM configurations are evaluated in uncalibrated mode. Green is best and light green is second best.
Method Sequence Avg
360 desk desk2 floor plant room rpy teddy xyz
Ours ($\mathrm{Sim}(3)$, $w=8$) 0.070 0.026 0.030 0.048 0.026 0.081 0.024 0.035 0.015 0.040
Ours ($\mathrm{Sim}(3)$, $w=16$) 0.112 0.045 0.123 0.261 0.022 0.137 0.044 0.044 0.016 0.089
Ours ($\mathrm{SL}(4)$, $w=8$) 0.179 0.046 0.095 0.210 0.033 0.272 0.038 0.130 0.031 0.115
Ours ($\mathrm{SL}(4)$, $w=16$) 0.147 0.032 0.087 0.158 0.027 0.150 0.037 0.069 0.035 0.083
Table 6: Dense reconstruction evaluation on 7-Scenes [58] (unit: m).
Method 7-Scenes
Uncalib. MASt3R-SLAM* [44] 0.066 0.068 0.045 0.056
Ours ($\mathrm{Sim}(3)$, $w=1$) 0.070 0.066 0.051 0.059
Ours ($\mathrm{Sim}(3)$, $w=8$) 0.067 0.054 0.056 0.055
Ours ($\mathrm{Sim}(3)$, $w=16$) 0.067 0.054 0.058 0.056
Ours ($\mathrm{Sim}(3)$, $w=32$) 0.067 0.052 0.062 0.057
Ours ($\mathrm{SL}(4)$, $w=1$) 0.090 0.080 0.068 0.074
Ours ($\mathrm{SL}(4)$, $w=8$) 0.084 0.067 0.065 0.066
Ours ($\mathrm{SL}(4)$, $w=16$) 0.075 0.061 0.063 0.060
Ours ($\mathrm{SL}(4)$, $w=32$) 0.067 0.052 0.058 0.055
B.2 Number of submaps per scene
As a reference, in Tables 7 and 8 we show the number of total submaps in each scene for 7-Scenes and TUM RGB-D for different values of experimented submap size, $w$, and also show the number of loop closures in each scene.
Table 7: Window size $w$ and corresponding submap and loop closure counts when ${w}_{\text{loop}}=1$, shown as “# of submaps (# of loops)”.
Window size, $w$ Sequences in 7-Scenes [63]
chess fire heads office pumpkin kitchen stairs
1 29 (11) 50 (46) 62 (49) 58 (55) 43 (37) 43 (38) 14 (12)
8 4 (0) 7 (4) 8 (3) 8 (4) 6 (0) 6 (2) 2 (0)
16 2 (0) 4 (1) 4 (2) 4 (2) 3 (0) 3 (1) 1 (0)
32 1 (0) 2 (0) 2 (0) 2 (0) 2 (0) 2 (0) 1 (0)
Table 8: Window size $w$ and corresponding submap and loop closure counts when ${w}_{\text{loop}}=1$, shown as “# of submaps (# of loops)”.
Window size, $w$ Sequences in TUM-RGB-D [63]
360 desk desk2 floor plant room rpy teddy xyz
1 168 (151) 54 (42) 98 (84) 99 (87) 102 (92) 186 (162) 95 (89) 146 (125) 56 (54)
8 21 (4) 7 (4) 13 (7) 13 (3) 13 (5) 24 (7) 12 (10) 19 (9) 7 (5)
16 11 (2) 4 (2) 7 (4) 7 (2) 7 (2) 12 (4) 6 (4) 10 (4) 4 (2)
32 6 (1) 2 (0) 4 (2) 4 (1) 4 (2) 6 (2) 3 (1) 5 (2) 2 (0)
Appendix C Extra Qualitative Results
C.1 Extra examples of $\mathrm{SL}(4)$ versus $\mathrm{Sim}(3)$
While we have mentioned that the $\mathrm{Sim}(3)$ version of VGGT-SLAM often provides high quality reconstructions, here we provide additional examples of cases where $\mathrm{Sim}(3)$ is not sufficient and $\mathrm{SL}(4)$ is necessary to achieve consistent alignment across submaps.
Figure 4: Example on a tabletop scene showing $\mathrm{Sim}(3)$ is unable to align the submaps while $\mathrm{SL}(4)$ is able to correct for projective ambiguity. Here $w=32$ and $\tau_{\text{disparity}}=50$.
Figure 5: Example on a tabletop scene showing $\mathrm{Sim}(3)$ is unable to align the submaps while $\mathrm{SL}(4)$ is able to correct for projective ambiguity. The true scene only has one banana, but the $\mathrm{Sim}(3)$ reconstruction shows a hallucination of two caused by misalignment. Camera pose estimates are colored by submap. Here $w=16$ and $\tau_{\text{disparity}}=50$.
Figure 6: Example on an outdoor scene with yellow bollards surrounding tanks showing $\mathrm{Sim}(3)$ is unable to align the submaps while $\mathrm{SL}(4)$ is able to correct for projective ambiguity. The true scene has single bollards spaced around the tanks while the $\mathrm{Sim}(3)$ scene hallucinates clusters of bollards due to misalignment. Here $w=16$ and $\tau_{\text{disparity}}=25$.
C.2 7-Scenes Qualitative Results
Here we provide additional visualizations of scene reconstructions from the 7-Scenes dataset experiments for VGGT-SLAM with $\mathrm{SL}(4)$. We use the default parameters from Sec. 5.
Figure 7: Visualization of reconstruction on 7-Scenes fire scene with 2 submaps. Camera pose estimates are colored by submap.
Figure 8: Visualization of reconstruction on 7-Scenes heads scene with 2 submaps. Camera pose estimates are colored by submap. Part of the scene in cropped for visual clarity.
C.3 TUM RGB-D Qualitative Results
Here we provide additional visualizations of scene reconstructions from the TUM RGB-D dataset experiments for VGGT-SLAM with $\mathrm{SL}(4)$. We use the default parameters from Sec. 5.
Figure 9: Visualization of reconstruction on TUM room scene with 6 submaps. Camera pose estimates are colored by submap. Part of the scene in cropped for visual clarity.
Figure 10: Visualization of reconstruction on TUM 360 scene with 6 submaps. Camera pose estimates are colored by submap. Part of the scene in cropped for visual clarity.
Figure 11: Visualization of reconstruction on TUM xyz scene with 2 submaps. Camera pose estimates are colored by submap.